OpenStudy (anonymous):

How to do elimination process 3x+3y=27 x-3y=-11

OpenStudy (anonymous):

is the second equation supposed to have a y term ??

OpenStudy (anonymous):

Yes i'm sorry it's x-3y= -11

OpenStudy (anonymous):

thats perfect, because all you have to do is add the equations... 3x + 3y = 27 x - 3y = -11 ------------add you will see that the 3y's cancel out....then just add and what do you get ?

OpenStudy (anonymous):

Omg thanks :3 I really appreciate it , I finally got it

OpenStudy (anonymous):

some of them are not that simple...sometimes you have to multiply one or both of the equations by a certain number to make a variable term cancel out.

OpenStudy (anonymous):

Yes , I have problems with the multiplying ... It's one of the most things I get confused in

OpenStudy (anonymous):

for instance... 2x + 4y = 8 x + 3y = 7 for this one, you would multiply the second equation by -2 so the x's will be eliminated then you will have.. 2x + 4y = 8 -2x - 6y = -14 (result of multiplying by -2) -------------add -2y = - 6 y = 3 then you would sub 3 in for y in either of the original equations to find x. do you understand a little better ?

OpenStudy (anonymous):

Kind of but not much, this elimination process confuses me a little . There's some that I get it really good but then I suddenly get confused and I loose everything i did

OpenStudy (anonymous):

then there is the tougher ones....like this.. 2x + 3y = 4 3x + 4y = 10 do you know how you would go about solving this one ?

OpenStudy (anonymous):

Do you have to multiply?

OpenStudy (anonymous):

yes...you have to multiply a number to both equations either multiply the first equation by -3 and the second by 2 to cancel out the x's or you can multiply the first equation by -4 and the second equation by 3 to cancel out the y's. Either way, you should eventually arrive at the same answer

OpenStudy (anonymous):

the main idea is to get a variable to cancel out by making one be negative and one positive, of course keeping the same coefficient

OpenStudy (anonymous):

But one question... How are you supposed to know what number to multiply by ?

OpenStudy (anonymous):

you multiply by the number that will make one variable be the opposite of the other one...does that make sense ?

OpenStudy (anonymous):

Oh so it will be let's suppose .. One positive and one negative , right ?

OpenStudy (anonymous):

yes...but they have to have the same coefficient in order to cancel out, just one has to be positive and one negative......-2x and 2x cancel out.....3y and -3y cancel out...understand ?

OpenStudy (anonymous):

and if they do not have the same coefficient, you have to make them have the same coefficient, just one has to be negative and one has to be positive

OpenStudy (anonymous):

Wow now it all makes sense , wow you're really good thank you so much :3

OpenStudy (anonymous):

if you had this... 3x + ... x + .... what would you multiply to make the x's cancel out

OpenStudy (anonymous):

you would multiply the second one by -3 and then the x's would cancel out. I hope I helped you understand a little better

OpenStudy (anonymous):

I understand now :)

OpenStudy (anonymous):

good to hear :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

sure thing :)